Comerpickles:

Circle 1 has center (−6, 2) and a radius of 8 cm. Circle 2 has center (−1, −4) and a radius 6 cm. What transformations can be applied to Circle 1 to prove that the circles are similar? Enter your answers in the boxes. Enter the scale factor as a fraction in simplest form. The circles are similar because the transformation rule ( , ) can be applied to Circle 1 and then dilate it using a scale factor of

Comerpickles:

@Shadow

Comerpickles:

@tigerlover

Comerpickles:

@umm

Comerpickles:

@563blackghost

tigerlover:

All the circles are similar. That is because you can transform one circle onto another by two similarity operations: translation and dilation (scale factor). For these two circles you have to translate the center of circle 1 to the center of circle 2, which will make that the two circles are concentric (have the same center). The two centers are (-6,2) and (-1,- 4). To move the center (-6, 2) to (-1,4) you have to shift it 5 units to the right and 6 units down: -6 + 5 = -1 2 - 6 = - 4. After this translation, you dilate the circle with smaller radius using a scale factor equal to the ratio of the bigger radius to the smaller radius: 8/6 = 4/3. Enter the scale factor as a fraction in simplest form. ----> 4/3 The circles are similar because the transformation rule (+5 , - 6) can be applied to Circle 1 and then dilate it using a scale factor of ​4/3.